# Thermodynamically consistent Cahn-Hilliard-Navier-Stokes equations using the metriplectic dynamics formalism

### 16 Feb 2024 v1

@@ 2.3 Metriplectic 4-bracket for the Gibbs-Navier-Stokes system
Now let us turn to our fourth and final step of the algorithm, construction of the metriplectic 4-bracket. To this end we review the formalism of Morrison & Updike (2023) in general terms and then apply it to the GNS system, which is a generalization of an example given in that work.
@@ 2.3.1 General metriplectic 4-bracket dynamics
The Metriplectic 4-bracket theory was introduced by Morrison & Updike (2023) to describe the dissipative dynamic. Let us briefly recall the metriplectic 4-bracket description in infinite dimensions. In this description, we consider the dynamics of classical field theories with multi-component fields
@@
$(F,G;K,N)=$ $\int\!\!d^{N}\!z\!\!\int\!\!d^{N}\!z^{\prime}\,\!\!\int\!\!d^{N}\!z^{\prime\prime}\!\!\int\!\!d^{N}\!z^{\prime\prime\prime}\,\hat{R}^{\alpha\beta\gamma\delta}(z,z^{\prime},z^{\prime\prime},z^{\prime\prime\prime})$
×$\hskip 28.45274pt\times\frac{\delta F}{\delta\chi^{\alpha}(z)}\frac{\delta G}{\delta\chi^{\beta}\left(z^{\prime}\right)}\frac{\delta K}{\delta\chi^{\gamma}\left(z^{\prime\prime}\right)}\frac{\delta N}{\delta\chi^{\delta}\left(z^{\prime\prime\prime}\right)}\,,$ (33)
where $\hat{R}^{\alpha\beta\gamma\delta}$ is a 4-tensor functional operator with coordinate form given by the following integral kernel:
$\hat{R}^{\alpha\beta\gamma\delta}(z,z^{\prime},z^{\prime\prime},z^{\prime\prime\prime})[\chi]=\hat{R}(\mathbf{d}\chi^{\alpha}(z),\mathbf{d}\chi^{\beta}(z^{\prime}),\mathbf{d}\chi^{\gamma}(z^{\prime\prime}),\mathbf{d}\chi^{\delta}(z^{\prime\prime}))[\chi(z)]\,,$ (34)
@@
The minimal metriplectic properties of metriplectic 4-brackets are reminiscent of the algebraic properties possessed by a curvature tensor. In fact, every Riemannian manifold naturally has a metriplectic 4-bracket, and $(S,H;S,H)$ provides a notion of sectional curvature (see Morrison & Updike, 2023).
@@
$\partial_{t}o=\left(o,H;S,H\right)$ $=\int\!\!d^{N}\!z\!\!\int\!\!d^{N}\!z^{\prime}\,\!\!\int\!\!d^{N}\!z^{\prime\prime}\!\!\int\!\!d^{N}\!z^{\prime\prime\prime}\,\hat{R}^{\alpha\beta\gamma\delta}(z,z^{\prime},z^{\prime\prime},z^{\prime\prime\prime})$ (41)
×$\hskip 28.45274pt\times\frac{\delta o}{\delta\chi^{\alpha}(z)}\frac{\delta H}{\delta\chi^{\beta}(z^{\prime})}\frac{\delta S}{\delta\chi^{\gamma}(z^{\prime\prime})}\frac{\delta H}{\delta\chi^{\delta}(z^{\prime\prime\prime})}\,.$
If we choose $o$ to be the Hamiltonian $H$, then $\partial_{t}H=\left(H,H;S,H\right)\equiv 0$ by the antisymmetry condition of (36). If we choose $o$ to be the entropy $S$, then $\partial_{t}S=\left(S,H;S,H\right)\geq 0$ by (40).
@@
$\partial_{t}{\chi}^{\alpha}(z)=\left(\chi^{\alpha},H;S,H\right)=\!\!\int\!\!d^{N}\!z^{\prime\prime}\,G^{\alpha\beta}(z,z^{\prime\prime})\frac{\delta S}{\delta\chi^{\beta}(z^{\prime\prime})}\,,$ (42)
@@
$G^{\alpha\gamma}(z,z^{\prime\prime})\coloneqq\int\!\!d^{N}\!z^{\prime}\,\!\!\int\!\!d^{N}\!z^{\prime\prime}\,R^{\alpha\beta\gamma\delta}(z,z^{\prime},z^{\prime\prime},z^{\prime\prime\prime})\frac{\delta H}{\delta\chi^{\beta}(z^{\prime})}\frac{\delta H}{\delta\chi^{\delta}(z^{\prime\prime\prime})}\,.$ (43)
@@ 2.3.2 General Kulkarni-Nomizu construction
$(\Sigma\wedge M)\left(dF,dK,dG,dN\right)$ $=\,\Sigma\left(dF,dG\right)M\left(dK,dN\right)-\Sigma\left(dF,dN\right)M\left(dK,dG\right)$
$+M\left(dF,dG\right)\Sigma\left(dK,dN\right)-M\left(dF,dN\right)\Sigma\left(dK,dG\right)\,.$ (44)
A finite-dimensional form of $\Sigma$ would be a symmetric contravariant 2-tensor, say $\sigma$, and this would give the term
$\sigma(df,dg)=\sigma^{ij}\frac{\partial f}{\partial z^{i}}\frac{\partial g}{\partial z^{j}}\,.$ (45)
@@
$\Sigma(dF,dG)=\int\!\!d^{N}\!z\!\!\int\!\!d^{N}\!z^{\prime}\,L_{ab}^{\alpha\beta}(z,z^{\prime})\,\mathcal{L}^{a}\frac{\delta F}{\delta\chi^{\alpha}(z)}\,\mathcal{L}^{\prime b}\frac{\delta G}{\delta\chi^{\beta}(z^{\prime})}\,.$ (48)
@@
×$\qquad\times\ \frac{\delta F}{\delta\chi^{\alpha}(z)}\frac{\delta G}{\delta\chi^{\beta}(z^{\prime})}\,\frac{\delta K}{\delta\chi^{\gamma}(z^{\prime\prime})}\frac{\delta N}{\delta\chi^{\delta}(z^{\prime\prime\prime})}\,,$ (49)
@@
$\Sigma(dF,dG)(z)\coloneqq\!\int\!\!d^{N}\!z^{\prime}\,\Sigma^{\alpha\beta}(z,z^{\prime})\frac{\delta F}{\delta\chi^{\alpha}(z)}\frac{\delta G}{\delta\chi^{\beta}(z^{\prime})}=A^{\alpha\beta}(z)\frac{\delta F}{\delta\chi^{\alpha}(z)}\frac{\delta G}{\delta\chi^{\beta}(z)}\,,$ (50)
which could follow from (46) if we added an additional argument to $\Sigma$. Then, with a corresponding form for $M$ the algebraic curvature symmetries would be induced in the integrand. This is the case for our present purposes, where we assume the specific K-N form given in Morrison & Updike (2023), viz.  where the 4-bracket is given by
$(F,K;G,N)=\int d^{N}\!z\,W\,(\Sigma\wedge M)\left(dF,dK,dG,dN\right)\,,$ (51)
@@ 2.3.3 Metriplectic 4-bracket for the GNS system
$(F,K;G,N)$ $=\int_{\Omega}\nabla F_{\psi^{\alpha}}\cdot B^{\alpha\beta}\cdot\nabla G_{{\psi^{\beta}}}\,K_{\psi^{\gamma}}A^{\gamma\delta}N_{{\psi^{\delta}}}-\nabla F_{\psi^{\alpha}}\cdot B^{\alpha\beta}\cdot\nabla N_{{\psi^{\beta}}}\,K_{\psi^{\gamma}}A^{\gamma\delta}G_{{\psi^{\delta}}}$
$+\nabla K_{\psi^{\alpha}}\cdot B^{\alpha\beta}\cdot\nabla N_{{\psi^{\beta}}}\,F_{\psi^{\gamma}}A^{\gamma\delta}G_{{\psi^{\delta}}}-\nabla K_{\psi^{\alpha}}\cdot B^{\alpha\beta}\cdot\nabla G_{{\psi^{\beta}}}\,F_{\psi^{\gamma}}A^{\gamma\delta}N_{{\psi^{\delta}}}.$ (55)
@@
$+\nabla H_{\psi^{\alpha}}\cdot B^{\alpha\beta}\cdot\nabla H_{{\psi_{\beta}}}\,S_{\psi^{\gamma}}A^{\gamma\delta}S_{{\psi^{\delta}}}-2\,\nabla H_{\psi^{\alpha}}\cdot B^{\alpha\beta}\cdot\nabla S_{{\psi_{\beta}}}\,S_{\psi^{\gamma}}A^{\gamma\delta}H_{{\psi^{\delta}}}\,.$
@@
${\Sigma}(dF,dG)=\nabla F_{{\mathbf{m}}}:\bar{\bar{\Lambda}}_{1}:\nabla G_{{\mathbf{m}}}+\nabla F_{{\sigma}}\cdot\bar{\Lambda}_{2}\cdot\nabla G_{{\sigma}}+\nabla{F}_{\tilde{c}}\cdot\bar{\Lambda}_{3}\cdot\nabla{G}_{\tilde{c}}\,,$ (59)
@@
$(F,K;G,N)$ $=\int_{\Omega}\frac{1}{T}\Big{[}\left[K_{\sigma}\nabla F_{\mathbf{m}}-F_{\sigma}\nabla K_{\mathbf{m}}\right]:\bar{\bar{\Lambda}}:\left[N_{\sigma}\nabla G_{\mathbf{m}}-G_{\sigma}\nabla N_{\mathbf{m}}\right]$
@@
$+\big{[}K_{\sigma}\nabla F_{\tilde{c}}-F_{\sigma}\nabla K_{\tilde{c}}\big{]}\cdot\bar{D}\cdot\big{[}N_{\sigma}\nabla G_{\tilde{c}}-G_{\sigma}\nabla N_{\tilde{c}}\big{]}\Big{]}\,.$ (61)
@@
$\partial_{t}\mathbf{v}$ =$=\{\mathbf{v},H\}+(\mathbf{v},H;S,H)=-\mathbf{v}\cdot\nabla\mathbf{v}-\nabla p/\rho+\frac{1}{\rho}\nabla\cdot(\bar{\bar{\Lambda}}:\nabla\mathbf{v})\,,$ (63)
$\partial_{t}\rho$ =$=\{\rho,H\}+(\rho,H;S,H)=-\mathbf{v}\cdot\nabla\rho-\rho\,\nabla\cdot\mathbf{v}\,,$ (64)
$\partial_{t}\tilde{c}$ =$=\{\tilde{c},H\}+(\tilde{c},H;S,H)=-\mathbf{v}\cdot\nabla\tilde{c}-\tilde{c}\,\nabla\cdot\mathbf{v}+\nabla\cdot(\bar{D}\cdot\nabla\mu)\,,$ (65)
$\partial_{t}\sigma$ =$=\{\sigma,H\}+(\sigma,H;S,H)=-\mathbf{v}\cdot\nabla\sigma-\sigma\,\nabla\cdot\mathbf{v}$
$\qquad+\nabla\cdot\left(\frac{\bar{\kappa}}{T}\cdot\nabla T\right)+\frac{1}{T^{2}}\nabla T\cdot\bar{\kappa}\cdot\nabla T+\frac{1}{T}\nabla\mathbf{v}:\bar{\bar{\Lambda}}:\nabla\mathbf{v}+\frac{1}{T}\nabla\mu\cdot\bar{D}\cdot\nabla\mu\,.$ (66)
@@
$=\int_{\Omega}\frac{1}{T}\left[\nabla\mathbf{v}:\bar{\bar{\Lambda}}:\nabla\mathbf{v}+\frac{1}{T}\nabla T\cdot\bar{\kappa}\cdot\nabla T+\nabla\mu\cdot\bar{D}\cdot\nabla\mu\right]\geq 0\,.$ (67)
@@ 2.4.1 GNS metriplectic 2-bracket
As noted above we are concerned with the metriplectic dynamics introduced in Morrison (1984a, b, 1986) (see also Morrison, 2009; Coquinot & Morrison, 2020), but we mention that other binary brackets for describing dissipation were presented over the years (e.g. Kaufman & Morrison, 1982; Kaufman, 1984; Morrison & Hazeltine, 1984; Grmela, 1984; Grmela & Öttinger, 1997; Beris & Edwards, 1994; Edwards, 1998). In addition we mention a recent alternative approach to multiphase fluids, one based on constrained variational principles, that is given in Eldred & Gay-Balmaz (2020). We refer the reader to Morrison & Updike (2023) for comparisons with other formulations and how they emerge from the metriplectic 4-bracket.
@@
$\dot{H}=\{{H},\mathcal{F}\}-({H},\mathcal{F})_{H}=\{{H},{H}\}+\mathcal{T}({H},{S})_{H}=0\,;$ (71)
@@
$\dot{S}=\{S,\mathcal{F}\}-(S,\mathcal{F})_{H}=-({S},{H})_{H}+\mathcal{T}({S},{S})_{H}=\mathcal{T}({S},{S})_{H}\geq 0\,,$ (72)
which follows because $\{S,\mathcal{F}\}\equiv 0$ and $({S},{H})_{H}\equiv 0$. As shown in Morrison & Updike (2023) the metriplectic 2-bracket emerges from the 4-bracket as follows:
@@
$=\int_{\Omega}\nabla F_{\psi^{\alpha}}\cdot B^{\alpha\beta}\cdot\nabla G_{{\psi^{\beta}}}\,H_{\psi^{\gamma}}A^{\gamma\delta}H_{{\psi^{\delta}}}-\nabla F_{\psi^{\alpha}}\cdot B^{\alpha\beta}\cdot\nabla H_{{\psi^{\beta}}}\,H_{\psi^{\gamma}}A^{\gamma\delta}G_{{\psi^{\delta}}}$
$+\nabla H_{\psi^{\alpha}}\cdot B^{\alpha\beta}\cdot\nabla H_{{\psi_{\beta}}}\,F_{\psi^{\gamma}}A^{\gamma\delta}G_{{\psi^{\delta}}}-\nabla H_{\psi^{\alpha}}\cdot B^{\alpha\beta}\cdot\nabla G_{{\psi_{\beta}}}\,F_{\psi^{\gamma}}A^{\gamma\delta}H_{{\psi^{\delta}}}\,,$ (74)
@@
Another 2-bracket can be obtained from (55) by choosing a convenient choice of variables; viz., instead of the variables $\psi$ of (135) in (74) we choose the following:
@@
where without loss of generality we set $A^{66}=1$ and $B^{\alpha\beta}=L_{\alpha\beta}$ for this case. In §§ 2.4.2 we will follow Coquinot & Morrison (2020) and show how the bracket of (77) fits into the framework of conventional nonequilibrium thermodynamics as, e.g., described in de Groot & Mazur (1962). In this way we will physically identify the meaning of $L_{\alpha\beta}$.

### 29 Feb 2024 v2

@@ 2.3 Metriplectic 4-bracket for the Gibbs-Navier-Stokes system
Now let us turn to our fourth and final step of the algorithm, construction of the metriplectic 4-bracket. To this end we review the formalism of Morrison and Updike (2023) in general terms and then apply it to the GNS system, which is a generalization of an example given in that work.
@@ 2.3.1 General metriplectic 4-bracket dynamics
The metriplectic 4-bracket theory was introduced by Morrison and Updike (2023) to describe the dissipative dynamic. Let us briefly recall the metriplectic 4-bracket description in infinite dimensions. In this description, we consider the dynamics of classical field theories with multi-component fields
@@
$(F,G;K,N)$ $=\int\!\!d^{N}\!z\!\!\int\!\!d^{N}\!z^{\prime}\,\!\!\int\!\!d^{N}\!z^{\prime\prime}\!\!\int\!\!d^{N}\!z^{\prime\prime\prime}\,\hat{R}^{\alpha\beta\gamma\delta}$
×$\times\frac{\delta F}{\delta\chi^{\alpha}(z)}\frac{\delta G}{\delta\chi^{\beta}\left(z^{\prime}\right)}\frac{\delta K}{\delta\chi^{\gamma}\left(z^{\prime\prime}\right)}\frac{\delta N}{\delta\chi^{\delta}\left(z^{\prime\prime\prime}\right)}\,,$ (33)
where $\hat{R}^{\alpha\beta\gamma\delta}(z,z^{\prime},z^{\prime\prime},z^{\prime\prime\prime})$ is a 4-tensor functional operator with coordinate form given by the following integral kernel:
$\hat{R}^{\alpha\beta\gamma\delta}(z,z^{\prime},z^{\prime\prime},z^{\prime\prime\prime})[\chi]$ $=$ (34)
$\hat{R}(\mathbf{d}\chi^{\alpha}(z),\mathbf{d}\chi^{\beta}(z^{\prime}),\mathbf{d}\chi^{\gamma}(z^{\prime\prime}),\mathbf{d}\chi^{\delta}(z^{\prime\prime}))[\chi(z)]\,,$
@@
The minimal metriplectic properties of metriplectic 4-brackets are reminiscent of the algebraic properties possessed by a curvature tensor. In fact, every Riemannian manifold naturally has a metriplectic 4-bracket, and $(S,H;S,H)$ provides a notion of sectional curvature (see Morrison and Updike, 2023).
@@
$\partial_{t}o$ $=\left(o,H;S,H\right)=\int\!\!d^{N}\!z\!\!\int\!\!d^{N}\!z^{\prime}\,\!\!\int\!\!d^{N}\!z^{\prime\prime}\!\!\int\!\!d^{N}\!z^{\prime\prime\prime}\,\hat{R}^{\alpha\beta\gamma\delta}$
×$\times\frac{\delta o}{\delta\chi^{\alpha}(z)}\frac{\delta H}{\delta\chi^{\beta}(z^{\prime})}\frac{\delta S}{\delta\chi^{\gamma}(z^{\prime\prime})}\frac{\delta H}{\delta\chi^{\delta}(z^{\prime\prime\prime})}\,.$ (41)
If we choose $o$ to be the Hamiltonian $H$, then $\dot{H}=\left(H,H;S,H\right)\equiv 0$ by the antisymmetry condition of (36). If we choose $o$ to be the entropy $S$, then $\dot{S}=\left(S,H;S,H\right)\geq 0$ by (40).
@@
$\partial_{t}{\chi}^{\alpha}(z)$ $=\left(\chi^{\alpha},H;S,H\right)$
$=\!\!\int\!\!d^{N}\!z^{\prime\prime}\,G^{\alpha\beta}(z,z^{\prime\prime})\frac{\delta S}{\delta\chi^{\beta}(z^{\prime\prime})}\,,$ (42)
@@
$G^{\alpha\gamma}(z,z^{\prime\prime})$ $\coloneqq\int\!\!d^{N}\!z^{\prime}\,\!\!\int\!\!d^{N}\!z^{\prime\prime}\,R^{\alpha\beta\gamma\delta}(z,z^{\prime},z^{\prime\prime},z^{\prime\prime\prime})$
×$\hskip 28.45274pt\times\frac{\delta H}{\delta\chi^{\beta}(z^{\prime})}\frac{\delta H}{\delta\chi^{\delta}(z^{\prime\prime\prime})}\,.$ (43)
@@ 2.3.2 General Kulkarni-Nomizu construction
$(\Sigma\wedge M)\left(dF,dK,dG,dN\right)$ $=\,\Sigma\left(dF,dG\right)M\left(dK,dN\right)$
$-\Sigma\left(dF,dN\right)M\left(dK,dG\right)$
$+M\left(dF,dG\right)\Sigma\left(dK,dN\right)$
$-M\left(dF,dN\right)\Sigma\left(dK,dG\right)\,.$ (44)
A finite-dimensional form of $\Sigma$ would be a symmetric contravariant 2-tensor, say $\gamma$, and this would give the term
$\gamma(df,dg)=\gamma^{ij}\frac{\partial f}{\partial z^{i}}\frac{\partial g}{\partial z^{j}}\,.$ (45)
@@
$\Sigma(dF,dG)$ $=\int\!\!d^{N}\!z\!\!\int\!\!d^{N}\!z^{\prime}\,L_{ab}^{\alpha\beta}(z,z^{\prime})$
×$\hskip 28.45274pt\times\mathcal{L}^{a}\frac{\delta F}{\delta\chi^{\alpha}(z)}\,\mathcal{L}^{\prime b}\frac{\delta G}{\delta\chi^{\beta}(z^{\prime})}\,.$ (48)
@@
×$\qquad\qquad\times\ \frac{\delta F}{\delta\chi^{\alpha}(z)}\frac{\delta G}{\delta\chi^{\beta}(z^{\prime})}\,\frac{\delta K}{\delta\chi^{\gamma}(z^{\prime\prime})}\frac{\delta N}{\delta\chi^{\delta}(z^{\prime\prime\prime})}\,,$ (49)
@@
$\Sigma(dF,dG)(z)$ $\coloneqq\!\int\!\!d^{N}\!z^{\prime}\,\Sigma^{\alpha\beta}(z,z^{\prime})\frac{\delta F}{\delta\chi^{\alpha}(z)}\frac{\delta G}{\delta\chi^{\beta}(z^{\prime})}$
$=A^{\alpha\beta}(z)\frac{\delta F}{\delta\chi^{\alpha}(z)}\frac{\delta G}{\delta\chi^{\beta}(z)}\,,$ (50)
which could follow from (46) if we added an additional argument to $\Sigma$. Then, with a corresponding form for $M$ the algebraic curvature symmetries would be induced in the integrand. This is the case for our present purposes, where we assume the specific K-N form given in Morrison and Updike (2023), viz.  where the 4-bracket is given by
$(F,K;G,N)=\!\!\!\int d^{N}\!z\,W\,(\Sigma\wedge M)\left(dF,dK,dG,dN\right)\,,$
@@ 2.3.3 Metriplectic 4-bracket for the GNS system
$(F,K;G,N)$ $=\int_{\Omega}\nabla F_{\psi^{\alpha}}\cdot B^{\alpha\beta}\cdot\nabla G_{{\psi^{\beta}}}\,K_{\psi^{\gamma}}A^{\gamma\delta}N_{{\psi^{\delta}}}$
$-\nabla F_{\psi^{\alpha}}\cdot B^{\alpha\beta}\cdot\nabla N_{{\psi^{\beta}}}\,K_{\psi^{\gamma}}A^{\gamma\delta}G_{{\psi^{\delta}}}$
$+\nabla K_{\psi^{\alpha}}\cdot B^{\alpha\beta}\cdot\nabla N_{{\psi^{\beta}}}\,F_{\psi^{\gamma}}A^{\gamma\delta}G_{{\psi^{\delta}}}$
$-\nabla K_{\psi^{\alpha}}\cdot B^{\alpha\beta}\cdot\nabla G_{{\psi^{\beta}}}\,F_{\psi^{\gamma}}A^{\gamma\delta}N_{{\psi^{\delta}}}\,.$ (54)
@@
$+\nabla H_{\psi^{\alpha}}\cdot B^{\alpha\beta}\cdot\nabla H_{{\psi_{\beta}}}\,S_{\psi^{\gamma}}A^{\gamma\delta}S_{{\psi^{\delta}}}$
$-2\,\nabla H_{\psi^{\alpha}}\cdot B^{\alpha\beta}\cdot\nabla S_{{\psi_{\beta}}}\,S_{\psi^{\gamma}}A^{\gamma\delta}H_{{\psi^{\delta}}}\,.$
@@
${\Sigma}(dF,dG)$ $=\nabla F_{{\mathbf{m}}}:\bar{\bar{\Lambda}}_{1}:\nabla G_{{\mathbf{m}}}+\nabla F_{{\sigma}}\cdot\bar{\Lambda}_{2}\cdot\nabla G_{{\sigma}}$
$\hskip 56.9055pt+\nabla{F}_{\tilde{c}}\cdot\bar{\Lambda}_{3}\cdot\nabla{G}_{\tilde{c}}\,,$ (58)
@@
$(F,K;G,N)$ $=$ (60)
$\int_{\Omega}\!\frac{1}{T}\Big{[}\left[K_{\sigma}\nabla F_{\mathbf{m}}-F_{\sigma}\nabla K_{\mathbf{m}}\right]:\bar{\bar{\Lambda}}:\left[N_{\sigma}\nabla G_{\mathbf{m}}-G_{\sigma}\nabla N_{\mathbf{m}}\right]$
@@
$+\big{[}K_{\sigma}\nabla F_{\tilde{c}}-F_{\sigma}\nabla K_{\tilde{c}}\big{]}\cdot\bar{D}\cdot\big{[}N_{\sigma}\nabla G_{\tilde{c}}-G_{\sigma}\nabla N_{\tilde{c}}\big{]}\Big{]}\,.$
@@
$\partial_{t}\mathbf{v}$ =$=\{\mathbf{v},H\}+(\mathbf{v},H;S,H)$
$=-\mathbf{v}\cdot\nabla\mathbf{v}-\nabla p/\rho+\frac{1}{\rho}\nabla\cdot(\bar{\bar{\Lambda}}:\nabla\mathbf{v})\,,$ (62)
$\partial_{t}\rho$ =$=\{\rho,H\}+(\rho,H;S,H)$
$=-\mathbf{v}\cdot\nabla\rho-\rho\,\nabla\cdot\mathbf{v}\,,$ (63)
$\partial_{t}\tilde{c}$ =$=\{\tilde{c},H\}+(\tilde{c},H;S,H)$
$=-\mathbf{v}\cdot\nabla\tilde{c}-\tilde{c}\,\nabla\cdot\mathbf{v}+\nabla\cdot(\bar{D}\cdot\nabla\mu)\,,$ (64)
$\partial_{t}\sigma$ =$=\{\sigma,H\}+(\sigma,H;S,H)$
$=-\mathbf{v}\cdot\nabla\sigma-\sigma\,\nabla\cdot\mathbf{v}$
$\qquad+\nabla\cdot\left(\frac{\bar{\kappa}}{T}\cdot\nabla T\right)+\frac{1}{T^{2}}\nabla T\cdot\bar{\kappa}\cdot\nabla T$
$\qquad+\frac{1}{T}\nabla\mathbf{v}:\bar{\bar{\Lambda}}:\nabla\mathbf{v}+\frac{1}{T}\nabla\mu\cdot\bar{D}\cdot\nabla\mu\,.$ (65)
@@
$=\int_{\Omega}\frac{1}{T}\bigg{[}\nabla\mathbf{v}:\bar{\bar{\Lambda}}:\nabla\mathbf{v}+\frac{1}{T}\nabla T\cdot\bar{\kappa}\cdot\nabla T$
$\hskip 88.2037pt+\nabla\mu\cdot\bar{D}\cdot\nabla\mu\bigg{]}\geq 0\,.$ (66)
@@ 2.4.1 GNS metriplectic 2-bracket
As noted above we are concerned with the metriplectic dynamics introduced in Morrison (1984a, b, 1986) (see also Morrison, 2009; Coquinot and Morrison, 2020), but we mention that other binary brackets for describing dissipation were presented over the years (e.g. Kaufman and Morrison, 1982; Kaufman, 1984; Morrison and Hazeltine, 1984; Grmela, 1984; Grmela and Öttinger, 1997; Beris and Edwards, 1994; Edwards, 1998). In addition we mention a recent alternative approach to multiphase fluids, one based on constrained variational principles, that is given in Eldred and Gay-Balmaz (2020). We refer the reader to Morrison and Updike (2023) for comparisons with other formulations and how they emerge from the metriplectic 4-bracket.
@@
$\dot{H}$ =$=\{{H},\mathcal{F}\}-({H},\mathcal{F})_{H}$
=$=\{{H},{H}\}+\mathcal{T}({H},{S})_{H}=0\,;$ (70)
@@
$\dot{S}$ =$=\{S,\mathcal{F}\}-(S,\mathcal{F})_{H}$
$=-({S},{H})_{H}+\mathcal{T}({S},{S})_{H}=\mathcal{T}({S},{S})_{H}\geq 0\,,$ (71)
which follows because $\{S,\mathcal{F}\}\equiv 0$ and $({S},{H})_{H}\equiv 0$. As shown in Morrison and Updike (2023) the metriplectic 2-bracket emerges from the 4-bracket as follows:
@@
$=\int_{\Omega}\nabla F_{\psi^{\alpha}}\cdot B^{\alpha\beta}\cdot\nabla G_{{\psi^{\beta}}}\,H_{\psi^{\gamma}}A^{\gamma\delta}H_{{\psi^{\delta}}}$
$-\nabla F_{\psi^{\alpha}}\cdot B^{\alpha\beta}\cdot\nabla H_{{\psi^{\beta}}}\,H_{\psi^{\gamma}}A^{\gamma\delta}G_{{\psi^{\delta}}}$
$+\nabla H_{\psi^{\alpha}}\cdot B^{\alpha\beta}\cdot\nabla H_{{\psi_{\beta}}}\,F_{\psi^{\gamma}}A^{\gamma\delta}G_{{\psi^{\delta}}}$
$-\nabla H_{\psi^{\alpha}}\cdot B^{\alpha\beta}\cdot\nabla G_{{\psi_{\beta}}}\,F_{\psi^{\gamma}}A^{\gamma\delta}H_{{\psi^{\delta}}}\,,$ (73)
@@
Another 2-bracket can be obtained from (54) by making a convenient choice of variables; viz., instead of the variables $\psi$ of (134) in (73) we choose the following:
@@
where without loss of generality we set $A^{66}=1$ and $B^{\alpha\beta}=L_{\alpha\beta}$ for this case. In §§ 2.4.2 we will follow Coquinot and Morrison (2020) and show how the bracket of (76) fits into the framework of conventional nonequilibrium thermodynamics as, e.g., described in de Groot and Mazur (1962). In this way we will physically identify the meaning of $L_{\alpha\beta}$.